OpenStudy (anonymous):

(((word problem)))) ---Winning percentage. In a basketball league, the Falcons won 15 of their first 20 games. In order to win 60% of the total number of games, how many more games will they have to play, assuming they win only half of the remaining games?

OpenStudy (a_clan):

(15+x)/(20+y)=(60/100).....(1) and x=(y/2) or y=2x...........(2) Substitute (2) in (1)

OpenStudy (gw2011):

They have to play 30 more games. Here is how I solved it: 15+0.50x and 20+x Where x=the number of additional games to be played (15+0.50x)/(20+x)=0.60 Multiply both sides by (20+x) to eliminate denominator 15+0.50x=(0.60)(20+x) 15+0.50x=12+0.60x 15-12=0.60x-0.50x 3=0.10x Rearranging this equation you get: 0.10x=3 x=30 This is the number of additional games to be played